OpenStudy (anonymous):

5 6th grade question!

OpenStudy (anonymous):

Find the area of a square tile that is 3 feet long on each side. A. 3 ft² B. 6 ft² C. 9 ft² D. 12 ft²

OpenStudy (anonymous):

Area of a square: \(A = s^2=3^2=?\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

NO!!!!!!!!!!!!!!!!!!!c

OpenStudy (anonymous):

yes c. 9 sq ft

OpenStudy (anonymous):

thanl you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Multiply. 40.631 · 0.17 A. 6.90727 B. 6.9727 C. 69.0727 D. 690.727

OpenStudy (anonymous):